If 2.40g of CuSO4 is dissolved in 9.95*10^2 ml of 0.350 M NH3, calculate the concentrations of the following species at equilibrium. (a) Cu2+ (b)NH3 (c)Cu(NH3)4^2+
1
Expert's answer
2020-02-04T06:34:24-0500

C0(NH3)=0.35M

n(CuSO4)=2.40/159.6=0.015 mol

C0(Cu2+)=0.015/0.995=0.0151 M


Cu2++4NH3=Cu(NH3)42+


k1234 = [Cu2+]*[NH3]4/[Cu(NH3)42+] = 9.33*10-13 (instability constant of this complex)

(0.0151-x)*(0.35-4x)4/x=9.33*10-13

x=0.0151 = [Cu(NH3)42+] (ammost all Cu2+ react with NH3)

[NH3] = 0.35-0.0151*4=0.2896M

(0.0151-x)*0.294/0.0151=9.33*10-13

(0.0151-x)=1.99*10-12 = [Cu2+]

Answer: [Cu2+]=1.99*10-12M; [NH3] =0.2896M; [Cu(NH3)42+]=0.0151M